1. Solve using the Laplace transform
y" − 6y' + 18y = 36 y(0) = 1, y'(0) = 6

3. Solve

t
f(t)−cos2t + ∫ f(τ)sin(t−τ)dτ =1

0
Answer #1

Isor Giren yl-6y! H8y-36 y(0)=1 vlo) We know defyll & Sy(s)-SyLo) - Y%) 0 daviye SYS) - Y0) Apply Paplace to ean 6 dly } – 6=) 3²+36= AS²+(-6A+B)S HISAIC - A-1, 6A+B=0/184+C = 36 =) B=64 ) 1811)+(=36 -) B=6 = 36 T 526SH8 S- S Apply laplare inverse d

Plz plz plz do Upvote